# Gradient-flow adaptive importance sampling for Bayesian leave one out cross-validation with application to sigmoidal classification models

### v2

@@ 2.1 Adaptive importance sampling
$\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}\left[f(\boldsymbol{\theta})\right]\approx\sum_{k=1}^{s}\frac{\chi_{ik}}{\sum_{j=1}^{s}\chi_{ij}}f(\boldsymbol{\phi}_{k})$ (14)
$\qquad\chi_{ik}=\frac{\mathcal{J}_{i}(\boldsymbol{\theta}_{k})}{\hat{\pi}(\boldsymbol{\theta}_{k}|\mathcal{D})}\pi(\boldsymbol{\phi}_{k})\prod_{j\neq i}\ell(\boldsymbol{\phi}_{k}|\mathbf{x}_{j},y_{j}),$ (15)
where $\pi(\boldsymbol{\phi}_{k})$ is the prior density at $\boldsymbol{\phi}_{k}$, canceling out the two unknown constants corresponding to $\pi(\boldsymbol{\phi}_{k}|\mathcal{D})$ and $\nu_{i}.$
@@ 3 Methods
$y_{i}|\boldsymbol{\theta},\mathbf{x}_{i}$ $\ {\sim}\ \textrm{Bernoulli}\left(\mathscr{p}_{i}(\boldsymbol{\theta})\right)$
$\ell(\boldsymbol{\theta}|y_{i},\mathbf{x}_{i})$ $=\mathscr{p}_{i}(\boldsymbol{\theta})^{y_{i}}(1-\mathscr{p}_{i}(\boldsymbol{\theta}))^{1-y_{i}},$ (16)
and where $\mathscr{p}_{i}(\boldsymbol{\theta})\equiv\mathscr{p}(\boldsymbol{\theta},\mathbf{x}_{i})$ is the predicted outcome probability for observation $i$, and computes the statistics of the parameter vector $\boldsymbol{\theta}$ by applying Bayes rule
$\pi(\boldsymbol{\theta}|\mathcal{D})\propto\pi(\boldsymbol{\theta})\prod_{i=1}^{n}\ell(\boldsymbol{\theta}|y_{i},\mathbf{x}_{i}),$ (17)
where $\pi(\boldsymbol{\theta})$ is the joint prior distribution over the parameter vector $\boldsymbol{\theta}$, corresponding to the likelihood function $\ell(\boldsymbol{\theta})$. There are usually many plausible choices for $\ell,$ even within the same family of models (for example different sets of features in logistic regression) – one needs a method to evaluate different models to inform decisions such as model selection.
Eq. 11 is valid for an arbitrary bijection $T_{i}.$ The objective of using transformations is to shift the proposal distribution closer to the targeted LOO distribution for each observation – to partially invert the Bayes rule. To this end, we motivate two different transformations of the form
$T_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}+hQ_{i}(\boldsymbol{\theta}),$ (18)
for some small step-size $h$ and function $Q_{i}.$
@@ 3.1 Gradient flow transformations
KL divergence descent: We consider choosing $T_{i}$ to minimize the KL divergence $\mathrm{D}_{\mathrm{KL}}\left(\pi({\boldsymbol{\theta}|\mathcal{D}^{(-i)}})\|\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\right)$, which is equivalent to minimizing the cross-entropy with respect to the mapping $T_{i}$,
$H\left(\pi({\boldsymbol{\theta}|\mathcal{D}^{(-i)}}),\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\right)=$
$\qquad-\int\nu_{i}(\boldsymbol{\phi})\pi(\boldsymbol{\phi}|\mathcal{D})\log\frac{\pi(T_{i}^{-1}(\boldsymbol{\phi})|\mathcal{D})}{\mathcal{J}_{T_{i}}(T_{i}^{-1}(\boldsymbol{\phi}))}\mathrm{d}\boldsymbol{\phi}.$ (19)
@@
$\frac{\partial T_{i}(\boldsymbol{\theta},t)}{\partial t}=-\frac{\delta H\left(\pi({\boldsymbol{\theta}|\mathcal{D}^{(-i)}}),\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\right)}{\delta T_{i}}$ (20)
and use this fact to refine, using the method of lines, an initial guess of $T_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}$ with forward Euler discretization of step-size $h[\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}[\ell(\boldsymbol{\theta},\mathbf{x}_{i},y_{i})]]^{-1}$, for $0<h\ll 1$, to arrive at the transformation
$T^{\textrm{KL}}_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}-h\frac{\delta H\left(\pi({\boldsymbol{\theta}|\mathcal{D}^{(-i)}}),\pi_{\boldsymbol{\phi}}(\boldsymbol{\theta}|\mathcal{D})\right)}{\mathbb{E}_{\boldsymbol{\theta}|\mathcal{D}^{(-i)}}[\ell(\boldsymbol{\theta},\mathbf{x}_{i},y_{i})]\delta T_{i}}\Bigg{|}_{T(\boldsymbol{\theta})=\boldsymbol{\theta}}$
$=\boldsymbol{\theta}+h\underbrace{\pi({\boldsymbol{\theta}}|\mathcal{D})\nabla\left(\frac{1}{\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})}\right)}_{Q_{i}^{\textrm{KL}}}.$ (21)
Variance descent: In importance sampling, the variance of the estimator is conditional on the target function for expectation. Since we are interested in computing the LOO predictive probability for each observation $i$, it is natural to consider minimizing the variance of the transformed importance sampler for the function $\mathscr{p}_{i}(\boldsymbol{\theta})=\mathscr{p}(\boldsymbol{\theta},\mathbf{x}_{i}).$ However, this objective yields a transformation that is only useful for observations where $y_{i}=0$ (see Supplemental MaterialsS.1.2). Instead, we seek to minimize the variance with respect to estimating the complement probability $\mathscr{p}_{i}(\boldsymbol{\theta})^{1-y_{i}}(1-\mathscr{p}_{i}(\boldsymbol{\theta}))^{y_{i}}.$
@@
$T^{\textrm{Var}}_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}+hQ_{i}^{\textrm{Var}}(\boldsymbol{\theta})$
$Q_{i}^{\textrm{Var}}(\boldsymbol{\theta})$ $=\pi(\boldsymbol{\theta}|\mathcal{D})\left[\frac{1-\mathscr{p}_{i}(\boldsymbol{\theta})}{\mathscr{p}_{i}(\boldsymbol{\theta})}\right]^{2y_{i}-1}$
×$\qquad\times\nabla\left(\left[\frac{1-\mathscr{p}_{i}(\boldsymbol{\theta})}{\mathscr{p}_{i}(\boldsymbol{\theta})}\right]^{2y_{i}-1}\right).$ (22)
Resolving the posterior density: Both the KL (Eq. 21) and variance (Eq. 22) descent transformations take steps proportional to the posterior density $\pi(\boldsymbol{\theta}|\mathcal{D}).$ If a variational approximation for $\pi({\boldsymbol{\theta}}|\mathcal{D})$ is available, using it in Eqs. 21 and 22 as a stand-in for the posterior density helps simplify the computation of the transformations and their Jacobians, particularly when using mean-field or low-order Automatic Differentiation Variational Inference (ADVI) [Kucukelbir et al., 2017, Blei et al., 2017].
In the absence of variational approximation, one may evaluate the posterior densities exactly using the Bayes rule, absorbing the unknown normalization constant $\mathcal{Z}$ into the step size $h.$ Then, the KL and variance descent transformations are
$T^{\textrm{KL}}_{i}(\boldsymbol{\theta})$
$=\boldsymbol{\theta}+h\pi(\boldsymbol{\theta})\left[\prod_{j}\ell(\boldsymbol{\theta}|\mathbf{x}_{j},y_{j})\right]\nabla\left(\frac{1}{\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i})}\right),$ (23)
and $T^{\textrm{Var}}_{i}(\boldsymbol{\theta})=\boldsymbol{\theta}+hQ_{i}^{\textrm{Var}}(\boldsymbol{\theta}),$ where
$Q_{i}^{\textrm{Var}}(\boldsymbol{\theta})$ $=\pi(\boldsymbol{\theta})\left[\prod_{j}\ell(\boldsymbol{\theta}|\mathbf{x}_{j},y_{j})\right]\left[\frac{1-\mathscr{p}_{i}(\boldsymbol{\theta})}{\mathscr{p}_{i}(\boldsymbol{\theta})}\right]^{2y_{i}-1}$
×$\qquad\times\nabla\left(\left[\frac{1-\mathscr{p}_{i}(\boldsymbol{\theta})}{\mathscr{p}_{i}(\boldsymbol{\theta})}\right]^{2y_{i}-1}\right).$ (24)
The obvious downside of using these exact transformations is the need to iterate over the entire dataset in order to evaluate the posterior density, which must be done for each parameter sample, for each data point.
@@
$\nabla\log[\mathcal{Z}\pi(\boldsymbol{\theta}|\mathcal{D})]=\nabla\log\pi(\boldsymbol{\theta})+\sum_{i}\nabla\log\ell(\boldsymbol{\theta}|\mathbf{x}_{i},y_{i}),$ (25)
@@
Step size selection The KL-divergence and variance descent transformations correspond to a forward Euler solver on the respective gradient flow equations. According to linear stability analysis, Euler’s method has the conditional stability criteria $h<{2}/{\max_{k}|\mathrm{Re}(\lambda_{k})|}$ where $\lambda_{k}$ are the eigenvalues of the Jacobian of the system (Jacobians of the functions $Q_{i}$). In each case the structure of the Jacobian admits cheap approximations of $\lambda_{k}$. However, for nonlinear systems, this criterion is not sufficient for achieving stability.
@@
$h_{i}=\rho\min_{s,\alpha}\left\{\left|{\sqrt{\Sigma_{\alpha,\alpha}}}/{Q_{i}(\boldsymbol{\theta}_{s})_{\alpha}}\right|\right\}$ (26)
where $\rho>0$ and $\sqrt{\Sigma_{\alpha,\alpha}}$ is the marginal posterior standard deviation of the $\alpha$-th component of $\boldsymbol{\theta}.$ This rule ensures that the transformation takes a step of at most $\rho$ posterior standard deviations in any parameter component. The objective of adaptation is to find any transformation that results in importance weights where the Pareto tail shape is sub-threshold. To this end, one can compute the transformations for a range of $\rho$ values in parallel using vectorized computations, saving computation at the cost of memory utilization.
Jacobian determinant approximation: For either single-step transformations, one may approximate $\left|\mathbf{J}_{T_{i}}\right|$ by noting that
$\mathcal{J}_{T_{i}}(\boldsymbol{\theta})=\left|1+h\nabla\cdot Q_{i}(\boldsymbol{\theta})\right|+\mathcal{O}(h^{2})$ (27)
and truncating to $\mathcal{O}(h),$ sidestepping the computation of Hessian matrices and their spectra. Note that any higher order terms in this expansion require characterization of the spectra of $\nabla Q_{i},$ for each observation $i$, and for each sampled parameter $\boldsymbol{\theta}_{k}$. For large problems, computing the Jacobian matrix and its spectra many times can become computationally problematic.
Overview: We have presented two gradient-flow descending transformations, each aimed at stabilizing a LOO importance sampler by bringing the proposal distribution closer to the LOO target in a different sense. The KL divergence and variance descent transformations correspond to the first-step of a forward Euler discretized solver for the corresponding gradient flow equations. The log-likelihood descent transformation repels the posterior parameter distribution away from the targeted LOO observation, undoing part of its contribution to the full data posterior. While each transformation uses gradient information, their Jacobians are simple to approximate, requiring no computation of full Hessian matrices.
Generally, one will find that many observations are amenable to direct importance sampling with $1/\ell$ weights (Eq. 3) in combination with Pareto smoothing (tail weight distribution shape parameter $\hat{k}<0.7$). One needs only transform the sampling distribution when the estimated shape parameter exceeds this threshold. For a given posterior sample of model parameters $\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{s}\overset{\mathrm{iid}}{\sim}\pi(\boldsymbol{\theta}|\mathcal{D})$, one undergoes the procedure for each given observation $i$:
@@
         Apply $T_{i}$ to each $\boldsymbol{\theta}_{k}$
@@
         if $\hat{k}\leq 0.7$ then
              Done
It is important to note that if any transformation takes $\hat{k}$ for a given observation under the threshold then adaptation is successful.
@@ 4 Examples
In this manuscript, we consider the broad widely-used class of models that have a sigmoidal parameterization
$\mathscr{p}_{i}(\boldsymbol{\theta})=p(\boldsymbol{\theta},\mathbf{x}_{i})=\sigma(\mu_{i}(\boldsymbol{\theta}))$ (28)
where $\sigma(\mu)=1/(1+e^{-\mu})$ is the sigmoid function and we denote $\mu_{i}(\boldsymbol{\theta})\equiv\mu(\boldsymbol{\theta},\mathbf{x}_{i})$ for some mean function $\mu$.
@@
$Q_{i}^{\textrm{KL}}(\boldsymbol{\theta})$ $=(-1)^{y_{i}}{\pi(\boldsymbol{\theta}|\mathcal{D})}e^{\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}\nabla\mu_{i}$ (29)
$Q_{i}^{\textrm{Var}}(\boldsymbol{\theta})$ $=(-1)^{y_{i}}\pi(\boldsymbol{\theta}|\mathcal{D})e^{2\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}\nabla\mu_{i},$ (30)
@@
$\mathbf{J}_{T_{i}^{\textrm{KL}}}(\boldsymbol{\theta})=\mathbf{I}+h(-1)^{y_{i}}{\pi(\boldsymbol{\theta}|\mathcal{D})}e^{\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}\Big{\{}\nabla\nabla\mu_{i}$
$\quad+\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+(1-2y_{i})\nabla\mu_{i}\right](\nabla\mu_{i})^{\intercal}\Big{\}}$ (32)
$\mathbf{J}_{T_{i}^{\textrm{Var}}}(\boldsymbol{\theta})=\mathbf{I}+h(-1)^{y_{i}}{\pi(\boldsymbol{\theta}|\mathcal{D})}e^{2\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}\Big{\{}\nabla\nabla\mu_{i}$
$\quad+\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+2(1-2y_{i})\nabla\mu_{i}\right](\nabla\mu_{i})^{\intercal}\Big{\}},$ (33)
@@
$\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})=\nabla\log\pi(\boldsymbol{\theta})$
$+\sum_{j}\left(y_{j}(1-\sigma(\mu_{j}))-(1-y_{j})\sigma(\mu_{j})\right)\nabla\mu_{j}(\boldsymbol{\theta}),$ (34)
and $\pi(\boldsymbol{\theta})$ is the prior. Here we will consider two popular sub-families of sigmoidal models.
Logistic Regression (LR): LR is a sigmoidal model where $\mu_{i}(\boldsymbol{\theta})=\mathbf{x}_{i}^{\intercal}{\boldsymbol{\beta}}$, So, $\nabla_{\boldsymbol{\beta}}\mu_{i}=\mathbf{x}_{i}$, and $\nabla\nabla\mu=\mathbf{0}$. Because the Hessian of $\mu$ vanishes, the Jacobian of the function $Q_{i}$ for each of the functions is a rank-one matrix and has only a single non-zero eigenvalue. LR admits exact Jacobian determinants for each of the transformations:
$\mathcal{J}_{T^{\textrm{KL}}_{i}}(\boldsymbol{\theta})=\Big{|}1+h(-1)^{y_{i}}{\pi(\boldsymbol{\theta}|\mathcal{D})}e^{\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}$
×$\quad\quad\quad\quad\quad\times\mathbf{x}_{i}^{\intercal}\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+(1-2y_{i})\mathbf{x}_{i}\right]\Big{|},$ (35)
$\mathcal{J}_{T_{i}^{\textrm{Var}}}(\boldsymbol{\theta})=\Big{|}1+h(-1)^{y_{i}}{\pi(\boldsymbol{\theta}|\mathcal{D})}e^{2\mu_{i}(\boldsymbol{\theta})(1-2y_{i})}$
×$\quad\quad\quad\quad\quad\times\mathbf{x}_{i}^{\intercal}\left[\nabla\log\pi(\boldsymbol{\theta}|\mathcal{D})+2(1-2y_{i})\mathbf{x}_{i}\right]\Big{|}.$ (36)
Bayesian (ReLU) Neural Networks: Bayesian ReLU-nets [Lee, 2000, Ghosh and Doshi-Velez, 2017, Choi et al., 2018, Kristiadi et al., 2020, Bhadra et al., 2019] are piecewise linear [Sudjianto et al., 2021, Wang, 2022, Montúfar et al., 2014, Sudjianto et al., 2020] extensions to regression models. Being locally linear, these models have block-sparse Hessians and are also amenable to some limited degree of interpretability [Sudjianto et al., 2020, Chang et al., 2023]. One may write an $L$-layer ReLU Bayesian neural network recursively

### v3

@@ 2.6 Correcting variational posteriors
$\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}\left[f(\bm{\theta})\right]\approx\sum_{k=1}^{s}\frac{\chi_{ik}}{\sum_{j=1}^{s}\chi_{ij}}f(\bm{\phi}_{k})\qquad\chi_{ik}=\frac{\mathcal{J}_{i}(\bm{\theta}_{k})}{\hat{\pi}(\bm{\theta}_{k}|\mathcal{D})}\pi(\bm{\phi}_{k})\prod_{j\neq i}\ell(\bm{\phi}_{k}|\mathbf{d}_{j}),$ (13)
where $\pi(\bm{\phi}_{k})$ is the prior density at $\bm{\phi}_{k}$, canceling out the two unknown constants corresponding to $\pi(\bm{\phi}_{k}|\mathcal{D})$ and $\nu_{i}.$
@@ 3 Methods
Eq. 10 is valid for an arbitrary bijection $T_{i}:\mathrm{supp}(\pi(\bm{\theta}))\to\mathrm{supp}(\pi(\bm{\theta})).$ The objective of using transformations is to shift the proposal distribution closer to the targeted LOO distribution for each observation – to invert the update version of Bayes’ rule. Returning to Eq. 1, in conjunction with Eq. 2, it is clear that the relative difference between the full posterior and the LOO posterior (for a given data point) is small – on the order of $1/N$. This fact motivates transformations of the form
$T_{i}(\bm{\theta})=\bm{\theta}+hQ_{i}(\bm{\theta}),$ (14)
for a perturbation parameter $h>0$ and function $Q_{i}.$
3.1 Partial moment-match stepping
The first two transformation methods presented in (Paananen et al., 2021) match the first moment and the first two moments respectively of the proposal distribution and the target distribution, independently for each scalar component of each model parameter. We generalize those transformations subject to a tunable scalar constant $\bar{h},$
$T_{\textrm{PMM}1}(\bm{\theta})$ $=\bm{\theta}+\bar{h}(\bar{\bm{\theta}}_{w}-\bar{\bm{\theta}})$ $T_{\textrm{PMM}2}(\bm{\theta})$ $=\bm{\theta}+\bar{h}\left(\mathbf{v}_{w}^{1/2}\circ\mathbf{v}^{-1/2}\circ(\bm{\theta}-\bar{\bm{\theta}})+\bar{\bm{\theta}}_{w}-\bm{\theta}\right)$
$\bar{\bm{\theta}}$ $=\frac{1}{s}\sum_{k=1}^{s}\bm{\theta}_{k}$ $\mathbf{v}$ $=\frac{1}{s}\sum_{k=1}^{s}(\bm{\theta}_{k}-\bar{\bm{\theta}})\circ(\bm{\theta}_{k}-\bar{\bm{\theta}})$
$\bar{\bm{\theta}}_{w}$ $=\frac{\sum_{k=1}^{s}\nu_{k}\bm{\theta}_{k}}{\sum_{k}\nu_{k}}$ $\mathbf{v}_{w}$ $=\frac{\sum_{k=1}^{s}\nu_{k}(\bm{\theta}_{k}-\bar{\bm{\theta}})\circ(\bm{\theta}_{k}-\bar{\bm{\theta}})}{\sum_{k=1}^{s}\nu_{k}}$ (15)
where setting $\bar{h}=1$ recovers the original transformations MM1/MM2 respectively.
@@ 3.2 Gradient flow transformations
3.2.1 KL divergence descent
We consider choosing $T_{i}$ to minimize the KL divergence $\mathrm{D}_{\mathrm{KL}}\left(\pi({\bm{\theta}|\mathcal{D}^{(-i)}})\|\pi_{\bm{\phi}}(\bm{\theta}|\mathcal{D})\right)$, which is equivalent to minimizing the cross-entropy with respect to the mapping $T_{i}$,
$H\left(\pi({\bm{\theta}|\mathcal{D}^{(-i)}}),\pi_{\bm{\phi}}(\bm{\theta}|\mathcal{D})\right)=-\int\nu_{i}(\bm{\phi})\pi(\bm{\phi}|\mathcal{D})\log\frac{\pi(T_{i}^{-1}(\bm{\phi})|\mathcal{D})}{\mathcal{J}_{T_{i}}(T_{i}^{-1}(\bm{\phi}))}\mathrm{d}\bm{\phi}.$ (16)
@@ 3.2.1 KL divergence descent
$\frac{\partial T_{i}(\bm{\theta},t)}{\partial t}=-\frac{\delta H\left(\pi({\bm{\theta}|\mathcal{D}^{(-i)}}),\pi_{\bm{\phi}}(\bm{\theta}|\mathcal{D})\right)}{\delta T_{i}}$ (17)
where $\delta/\delta T_{i}$ denotes the functional derivative of $H$ with respect to the transformation $T_{i},$ and use this fact to refine, using the method of lines, an initial guess of $T_{i}(\bm{\theta})=\bm{\theta}$ with forward Euler discretization of step size $h[\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}[\ell(\bm{\theta},\mathbf{x}_{i},y_{i})]]^{-1}$, for $0<h\ll 1$, to arrive at the transformation
$T^{\textrm{KL}}_{i}(\bm{\theta})=\bm{\theta}-\frac{h}{\mathbb{E}_{\bm{\theta}|\mathcal{D}^{(-i)}}[\ell(\bm{\theta}|\mathbf{d}_{i})]}\frac{\delta H\left(\pi({\bm{\theta}|\mathcal{D}^{(-i)}}),\pi_{\bm{\phi}}(\bm{\theta}|\mathcal{D})\right)}{\delta T_{i}}\Bigg{|}_{T(\bm{\theta})=\bm{\theta}}=\bm{\theta}+h\underbrace{\pi({\bm{\theta}}|\mathcal{D})\nabla\left(\frac{1}{\ell(\bm{\theta}|\mathbf{d}_{i})}\right)}_{Q_{i}^{\textrm{KL}}}$ (18)
3.2.2 Variance descent
In importance sampling, the variance of the estimator is conditional on the target function for expectation. Since we are interested in computing the LOO predictive probability for each observation $i$, it is natural to consider minimizing the variance of the transformed importance sampler for the function $\mathscr{p}_{i}(\bm{\theta})=\mathscr{p}(\bm{\theta},\mathbf{x}_{i}).$ However, this objective yields a transformation that is only useful for observations where $y_{i}=0$ (see Supplemental MaterialsS.1.2). Instead, we seek to minimize the variance with respect to estimating the complement probability $\mathscr{p}_{i}(\bm{\theta})^{1-y_{i}}(1-\mathscr{p}_{i}(\bm{\theta}))^{y_{i}}.$
@@ 3.2.2 Variance descent
$T^{\textrm{Var}}_{i}(\bm{\theta})=\bm{\theta}+hQ_{i}^{\textrm{Var}}(\bm{\theta})\qquad Q_{i}^{\textrm{Var}}(\bm{\theta})=\pi(\bm{\theta}|\mathcal{D})\frac{f(\bm{\theta})}{\ell(\bm{\theta}|\mathbf{d}_{i})}\nabla\left(\frac{f(\bm{\theta})}{\ell(\bm{\theta}|\mathbf{d}_{i})}\right)$ (19)
3.3 Resolving the posterior density
Both the KL (Eq. 18) and variance (Eq. 19) descent transformations take steps proportional to the posterior density $\pi(\bm{\theta}|\mathcal{D}).$ If a variational approximation for $\pi({\bm{\theta}}|\mathcal{D})$ is available, using it in Eqs. 18 and 19 as a stand-in for the posterior density helps simplify the computation of the transformations and their Jacobians, particularly when using mean-field or low-order Automatic Differentiation Variational Inference (ADVI) (Kucukelbir et al., 2017; Blei et al., 2017).
In the absence of variational approximation, one may evaluate the posterior densities exactly using the Bayes rule, absorbing the unknown normalization constant $\mathcal{Z}$ into the step size $h.$ The obvious downside of using these exact transformations is the need to iterate over the entire dataset in order to evaluate the posterior density, which must be done for each parameter sample, for each data point.
@@ 3.3 Resolving the posterior density
$\nabla\log[\mathcal{Z}\pi(\bm{\theta}|\mathcal{D})]=\nabla\log\pi(\bm{\theta})+\sum_{i}\nabla\log\ell(\bm{\theta}|\mathbf{d}_{i}),$ (20)
@@
3.4 Step size selection
The KL-divergence and variance descent transformations correspond to a forward Euler solver on the respective gradient flow equations. According to linear stability analysis, Euler’s method has the conditional stability criteria $h<{2}/{\max_{k}|\mathrm{Re}(\lambda_{k})|}$ where $\lambda_{k}$ are the eigenvalues of the Jacobian of the system (Jacobians of the functions $Q_{i}$). In each case the structure of the Jacobian admits inexpensive approximations of $\lambda_{k}$. However, for nonlinear systems, this criterion is not sufficient for achieving stability.
@@ 3.4 Step size selection
$h_{i}=\bar{h}\min_{s,\alpha}\left\{\left|\frac{\sqrt{\Sigma_{\alpha,\alpha}}}{Q_{i}(\bm{\theta}_{s})_{\alpha}}\right|\right\}$ (21)
where $\bar{h}>0$ and $\sqrt{\Sigma_{\alpha,\alpha}}$ is the marginal posterior standard deviation of the $\alpha$-th component of $\bm{\theta}.$ This rule ensures that the transformation takes a step of at most $\bar{h}$ posterior standard deviations in any parameter component. The objective of adaptation is to find any transformation that results in importance weights where the Pareto tail shape is sub-threshold. To this end, one can compute the transformations for a range of $\bar{h}$ values in parallel using vectorized computations, saving computation at the cost of memory utilization.
3.5 Jacobian determinant approximation
For either single-step transformations, one may approximate $\left|\mathbf{J}_{T_{i}}\right|$ by noting that
$\mathcal{J}_{T_{i}}(\bm{\theta})=\left|1+h\nabla\cdot Q_{i}(\bm{\theta})\right|+\mathcal{O}(h^{2})$ (22)
and truncating to $\mathcal{O}(h),$ sidestepping the computation of Hessian matrices and their spectra. Note that any higher order terms in this expansion require characterization of the spectra of $\nabla Q_{i},$ for each observation $i$, and for each sampled parameter $\bm{\theta}_{k}$. For large problems, computing the Jacobian matrix and its spectra many times can become computationally problematic.
3.6 Overview
We have presented four transformations, each aimed at stabilizing a LOO importance sampler by bringing the proposal distribution closer to the LOO target in a different sense. The PMM1/PMM2 transformations shift the moment of the posterior samples closer to that of the target distribution. The KL/Var descent transformations take one step along their corresponding gradient flow equations. While the latter two transformations use gradient information, their Jacobians are simple to approximate, requiring no computation of full Hessian matrices.
Generally, one will find that many observations are amenable to direct importance sampling with $1/\ell$ weights (Eq. 4) in combination with Pareto smoothing (tail weight distribution shape parameter $\hat{k}<0.7$). One needs only transform the sampling distribution when the estimated shape parameter exceeds this threshold. For a given posterior sample of model parameters $\bm{\theta}_{1},\ldots,\bm{\theta}_{s}\overset{\mathrm{iid}}{\sim}\pi(\bm{\theta}|\mathcal{D})$, one undergoes for each observation $i$ the following algorithm:
@@ 3.6 Overview
         Apply $T_{i}$ to each $\bm{\theta}_{k}$
@@
         if $\hat{k}\leq 0.7$ then Done
It is important to reiterate that if any transformation takes $\hat{k}$ for a given observation under the threshold then adaptation is successful – one avoids refitting the entire model.
@@ 4 Examples
$y_{i}|\bm{\theta},\mathbf{x}_{i}\ {\sim}\ \textrm{Bernoulli}\left(\mathscr{p}_{i}(\bm{\theta})\right)\qquad\ell(\bm{\theta}|y_{i},\mathbf{x}_{i})=\mathscr{p}_{i}(\bm{\theta})^{y_{i}}(1-\mathscr{p}_{i}(\bm{\theta}))^{1-y_{i}},$ (23)
and where $\mathscr{p}_{i}(\bm{\theta})\equiv\mathscr{p}(\bm{\theta},\mathbf{x}_{i})$ is the predicted outcome probability for observation $i$. In this manuscript, we pay special attention to the broad widely-used class of models that have a sigmoidal parameterization.
$\mathscr{p}_{i}(\bm{\theta})=p(\bm{\theta},\mathbf{x}_{i})=\sigma(\mu_{i}(\bm{\theta}))$ (24)
where $\sigma(\mu)=1/(1+e^{-\mu})$ is the sigmoid function and we denote $\mu_{i}(\bm{\theta})\equiv\mu(\bm{\theta},\mathbf{x}_{i})$ for some mean function $\mu$.
@@
$Q_{i}^{\textrm{KL}}(\bm{\theta})=(-1)^{y_{i}}{\pi(\bm{\theta}|\mathcal{D})}e^{\mu_{i}(\bm{\theta})(1-2y_{i})}\nabla\mu_{i}\qquad Q_{i}^{\textrm{Var}}(\bm{\theta})=(-1)^{y_{i}}\pi(\bm{\theta}|\mathcal{D})e^{2\mu_{i}(\bm{\theta})(1-2y_{i})}\nabla\mu_{i},$ (25)
@@
$\mathbf{J}_{T_{i}^{\textrm{KL/Var}}}(\bm{\theta})=\mathbf{I}+\Bigg{\{}h(-1)^{y_{i}}{\pi(\bm{\theta}|\mathcal{D})}e^{(1+1_{\textrm{Var}})\mu_{i}(\bm{\theta})(1-2y_{i})}$
×$\qquad\times\Big{\{}\nabla\nabla\mu_{i}+\left[\nabla\log\pi(\bm{\theta}|\mathcal{D})+(1+1_{\textrm{Var}})(1-2y_{i})\nabla\mu_{i}\right](\nabla\mu_{i})^{\intercal}\Big{\}}\Bigg{\}}$ (26)
@@
$\nabla\log\pi(\bm{\theta}|\mathcal{D})=\nabla\log\pi(\bm{\theta})+\sum_{j}\left(y_{j}(1-\sigma(\mu_{j}))-(1-y_{j})\sigma(\mu_{j})\right)\nabla\mu_{j}(\bm{\theta}),$ (27)
and $\pi(\bm{\theta})$ is the prior. Here we will consider two popular subfamilies of sigmoidal models.
4.1 Logistic Regression (LR)
LR is a sigmoidal model where $\mu_{i}(\bm{\theta})=\mathbf{x}_{i}^{\intercal}{\bm{\beta}}$, So, $\nabla_{\bm{\beta}}\mu_{i}=\mathbf{x}_{i}$, and $\nabla\nabla\mu=\mathbf{0}$. Because the Hessian of $\mu$ vanishes, the Jacobian of the function $Q_{i}$ for each of the functions is a rank-one matrix and has only a single non-zero eigenvalue. LR admits exact Jacobian determinants for each of the transformations:
$\mathcal{J}_{T^{\textrm{KL/Var}}_{i}}(\bm{\theta})=\Big{|}1+h(-1)^{y_{i}}{\pi(\bm{\theta}|\mathcal{D})}e^{(1+1_{\textrm{Var}})\mu_{i}(\bm{\theta})(1-2y_{i})}\mathbf{x}_{i}^{\intercal}\left[\nabla\log\pi(\bm{\theta}|\mathcal{D})+(1+1_{\textrm{Var}})(1-2y_{i})\mathbf{x}_{i}\right]\Big{|}.$ (28)
4.2 Bayesian (ReLU) Neural Networks
Bayesian ReLU-nets (Lee, 2000; Ghosh & Doshi-Velez, 2017; Choi et al., 2018; Kristiadi et al., 2020; Bhadra et al., 2019) are piecewise linear (Sudjianto et al., 2021; Wang, 2022; Montúfar et al., 2014; Sudjianto et al., 2020) extensions to regression models. Being locally linear, these models have block-sparse Hessians and are also amenable to some limited degree of interpretability (Sudjianto et al., 2020; Chang et al., 2023). One may write an $L$-layer ReLU Bayesian neural network recursively